OpenStudy (anonymous):

The faces of a tetrahedron die are numbered 1 to 4. If each of the 4 sides is equally likely to land down, what is the probability that the 2 does not end up face down?

OpenStudy (anonymous):

\(\frac 34 \) ?

OpenStudy (anonymous):

well, the chance for ANY number to land is 1/4 - (landing side)/(total sides) the chance that 2 will be the number that WON'T land, that is only 1,2, and 4 can land, means it will be 3/4 because the amount of sides that could land are 3, and the total sides are 4 thus 3/4

OpenStudy (anonymous):

that is only 1,3,4* can land lol woops

OpenStudy (anonymous):

The probability of 2 end up face down is \( \frac 14 \). Hnece probability of 2 not landing face down is \( 1 - \frac 14 =\frac 34 \).

OpenStudy (anonymous):

Hence*